OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360°. answers (2 square root of 2, 225°), (-2 square root of 2, 45°) (2 square root of 2, 135°), (-2 square root of 2, 315°) (2 square root of 2, 315°), (-2 square root of 2, 135°) (2 square root of 2, 45°), (-2 square root of 2, 225°)

OpenStudy (jdoe0001):

sit has been a bit laggy.. so

OpenStudy (anonymous):

what? @jdoe0001

OpenStudy (jdoe0001):

wooops I meant, the site gets laggy

OpenStudy (jdoe0001):

as in the website

OpenStudy (jdoe0001):

\(\large { \begin{array}{cccllll} (&2, &-2)\\ &\uparrow &\uparrow \\ &x&y \end{array}\qquad \begin{array}{llll} r^2=x^2+y^2\to {\color{brown}{ r}}=\sqrt{x^2+y^2} \\ \quad \\ tan(\theta)=\frac{y}{x}\to {\color{brown}{ \theta}}=tan^{-1}\left(\frac{y}{x}\right) \end{array} }\) what would that give you for the "r" or radius of those coordinates and for the angle?

OpenStudy (anonymous):

would the answer be (2 square root of 2, 45°), (-2 square root of 2, 225°) @jdoe0001

OpenStudy (jdoe0001):

hmmm what did you get for "r"?

OpenStudy (anonymous):

\[\sqrt{8}\]

OpenStudy (jdoe0001):

and... for \(\theta?\)

OpenStudy (anonymous):

or (2 square root of 2, 225°), (-2 square root of 2, 45°)

OpenStudy (anonymous):

for \[\theta \] 45

OpenStudy (jdoe0001):

well.... the inverse tangent function, has a range of \(\large -\frac{\pi}{2}<\theta<\frac{\pi}{2}\) so the 45 degree is is just a reference angle keep in mind that tangent of that angle, is a negative -1 since y/x = -1 and tangent is negative in only 2 places.... 2nd and 4th Quadrants because in the 2nd Quadrant, x is negative and y positive and in the 4th Quadrant, x is positive but y is negative